OpenStudy (anonymous):

root:3:(5)*root:3:(25) multiply to solve.

OpenStudy (anonymous):

\[\sqrt[3]{5} * \sqrt[3]{25} = \sqrt[3]{125} = 5 \] roots with the same ratial can be multiplied*

OpenStudy (anonymous):

Thank you that is what I got:)

OpenStudy (anonymous):

good ^^